Determine the molar enthalpy of the balanced reaction if 4441.9 kJ of energy is released when 100 g of C8H18 combusts with 400 g of O2.



1
Expert's answer
2021-10-15T02:55:04-0400

First write the equation,


C8H18"+\frac{25}{2}" O2"\longrightarrow" 8CO2+9H2O

From this equation we realize that 114g of C8H18 needs 400g of O2. There is 400g of oxygen but only 100g of Octane, implying that octane is the limiting reagent.

The following relation will be therefore applied.

100g Octane "\equiv" 4441.9KJ

Use the above relation find the joules that will be produced for one molar octane which is 114g.

Molar enthalpy= "\frac{114g/mol}{100g}×-4441.9KJ=-5063.766KJ/mol"

Note that the negative sign is used to denote that it's an exothermic reaction.